# A comprehensive survey on deep active learning and its applications in medical image analysis

### v1

@@ 3.3 Sampling Strategy
Most deep AL works used top-k to select samples with the highest informativeness for annotation. However, existing informativeness metrics face several issues, such as redundancy and class imbalance in selected samples. Instead of improving informativeness, we can introduce simple sampling strategies to resolve these issues effectively. Besides, specific sampling strategies can also be used for combining multiple informativeness metrics. Furthermore, with the recent development of deep AL, more studies directly employ neural networks for sample selection. In this context, we no longer evaluate informativeness but directly choose valuable samples from the unlabeled pool with neural networks. In summary, sampling strategies are crucial in AL, but prior surveys have seldom discussed their specific attributes. As one of the contributions of this survey, we systematically summarize different sampling strategies in AL, including diversity sampling, class-balanced sampling, hybrid sampling, and learnable sampling. The taxonomy of different sampling strategies in AL is shown in Fig. 5.
@@ 3.3.1 Diversity Sampling
Diversity strategies aim to reduce redundancy in selected samples. Sampling redundancy is a common issue for uncertainty-based and representativeness-based methods, meaning some selected samples are highly similar. The lack of diversity leads to the waste of the annotation budget. Besides, redundancy in the training set may cause deep models to overfit, thus degrading performance. Therefore, many AL methods employ diversity sampling to mitigate the redundancy in selected samples. In this section, we discuss four strategies of diversity sampling, including clustering, farthest-first traversal, determinantal point process (DPP), and specific strategies tailored to certain informativeness metrics.
@@
Clustering is one of the most commonly used strategies of diversity sampling. It groups the data into several clusters and then queries samples within each cluster. This strategy improves the coverage of the entire feature space, thereby easily boosting diversity. Ash et al. (2020) employed k-Means++ clustering on gradient embeddings to select diverse uncertain samples. Citovsky et al. (2021) boosted margin-based uncertainty sampling with hierarchical clustering. They selected samples with the smallest margins within each cluster. When the number of queries exceeded the number of clusters, samples from smaller clusters were prioritized. This method can extend to a huge annotation budget (e.g., one million). Jin et al. (2022a) employed BIRCH clustering and chose the samples with maximum information density within each cluster for labeling. Compared to k-Means, BIRCH clustering is less sensitive to outliers and can further identify noisy samples. In connectomics, Lin et al. (2020) trained two feature extractors with labeled and unlabeled samples, respectively. They then selected samples for annotation through multiple rounds of clustering. This method achieved excellent performance in synapse detection and mitochondria segmentation.
Farthest-first Traversal is also a widely used strategy for diverse queries. The farthest-first traversal requires the distance between sampling points to be as large as possible in the feature space. This leads to a more uniform distribution of selected samples within the feature space, thus improving the diversity of the sampling results. This technique was first adopted by Sener and Savarese (2018). Agarwal et al. (2020) and Caramalau et al. (2021) improved the diversity with farthest-first traversal, leveraging their proposed contextual diversity and GNN-augmented features, respectively. However, when the annotation budget is limited, the farthest-first traversal may be biased toward outliers.
@@
Determinantal point process is a stochastic probability model for selecting subsets from a larger set. DPP reduces the probability of sampling similar elements to ensure diversity in the results. Bıyık et al. (2019) employed two DPPs for sample selection: Uncertainty DPP is based on uncertainty scores, while Exploration DPP aims to find samples near decision boundaries. Then, sampling results from both DPPs were sent for expert annotation. However, DPP is more computationally intensive compared to clustering. Ash et al. (2020) compared the performance and time cost of using k-Means++ and k-DPP. Results showed that their performance is similar, but the time cost for k-Means++ is significantly lower than that for k-DPP.
Specific strategies: There are also specific strategies tailored to certain informativeness metrics. In uncertainty-based AL, BatchBALD (Kirsch et al., 2019) extended BALD-based uncertainty AL to batch mode. Results showed that BatchBALD improved the sampling diversity compared to (Gal et al., 2017). FI-based methods formulated AL as a semi-definite programming (SDP) problem to improve sampling diversity. Different methods were employed for solving SDP. Sourati et al. (2019) used a commercial solver to solve SDP, while Ash et al. (2021) proposed a greedy algorithm to adapt to high-dimensional feature space. Moreover, diversity is an essential part of representativeness-based AL. Cover-based AL inherently incorporates the considerations of diversity in its formulations. In discrepancy-based AL, Wasserstein distance is used for diverse query results (Shui et al., 2020; Mahmood et al., 2022). Density-based methods often employ strategies like clustering to improve diversity.
@@ 3.3.2 Class-balance Sampling
Class imbalance is a common issue for DL, where a small set of classes have many samples while the others only contain a few samples (Zhang et al., 2023). For example, in tasks such as medical image classification, normal samples often outnumber abnormal ones. Training on imbalanced datasets can lead to the overfitting of the majority classes and underfitting of the minority classes. Apart from dealing with class imbalance during training, AL mitigates class imbalance by avoiding over-annotation of the majority classes and enhancing the annotation of the minority classes during dataset construction.
@@
Classification: Choi et al. (2021b) directly estimated the probability of a classifier making a mistake for a given sample and decomposed it into three terms using Bayesian rules. First, they trained a VAE to estimate the likelihood of the data given a predicted class. Then, an additional classifier was trained upon VAE features to estimate class prior probabilities and the probability of mislabeling a specific class. By considering all three probabilities, they successfully mitigated class imbalance in AL. The proposed method achieved good performance on stepwise class-imbalanced CIFAR-10 and CIFAR-100 datasets. For uncertainty-based methods, Bengar et al. (2022) introduced an optimization framework to maintain class balance. They compensated the query of minority classes with the most confident samples of that class, leading to a more balanced class distribution in the queried dataset. In classification tasks, Munjal et al. (2022) tested various AL baselines on the long-tailed CIFAR-100 dataset. Results showed that no single method outperforms others on all budgets in the class-imbalance setting. However, as the number of labeled data increases, the performance gap between random sampling and the best AL method decreases. Hacohen et al. (2022) conducted experiments in a class-imbalanced setting similar to Munjal et al. (2022). The proposed TypiClust ensured class balance and outperformed other baseline methods. Jin et al. (2022c) assumed that samples closer to the tail of the distribution are more likely to belong to the minority classes. Thus, the tail probability is equivalent to the likelihood of minority classes. Specifically, they trained a VAE for feature extraction and adopted copula to estimate the tail probabilities upon VAE features. Finally, informative samples were selected with clustering and unequal probability sampling. The proposed method was validated on the ISIC 2020 dataset, which has a long-tailed distribution. Kothawade et al. (2022c) used submodular mutual information to focus more on samples of minority classes. They achieved excellent results on medical classification datasets in five different modalities, including X-rays, pathology, and dermoscopy. Besides, in blood cell detection under microscopy, Sadafi et al. (2019) requested expert annotation of a sample whenever its classification probability of the minority class exceeded 0.2.
@@
Segmentation: Due to some AL methods selecting regions instead of the entire image for annotation, there is a need to ensure that the selected regions contain rare or small objects (e.g., pedestrians or utility poles in autonomous driving). Cai et al. (2021) and Wu et al. (2022b) both proposed class-balanced sampling strategies for such scenarios, as detailed in §4.4.
@@ 3.3.3 Hybrid Sampling
In AL, some works may use multiple informativeness metrics simultaneously. Therefore, the effective integration of multiple metrics remains a critical issue. This issue is addressed by the hybrid sampling strategy discussed in this section. Two approaches to hybrid sampling are primarily used, including multi-round sampling and metric fusion.
Multi-round sampling first selects a subset of samples based on one particular informativeness metric and continues sample selection within this subset based on another informativeness metric. For example, SA (Yang et al., 2017) performed representativeness sampling based on uncertainty to reduce redundancy in the sampled set. Xie et al. (2022b) first selected samples with density-based methods and selected the most uncertain samples within each cluster of representative samples. In another study, Xie et al. (2022c) introduced distribution and data uncertainty based on EDL, and then a two-stage strategy was used for sample selection. Wu et al. (2022b) employed a more complex strategy, which sets dynamic weights to adjust the budget of representativeness and uncertainty sampling. The weight of representativeness sampling is larger initially, while the situation is reversed in the latter phase. This is because representativeness methods can quickly spot typical data, while uncertainty methods continuously improve the model by querying samples with erroneous predictions.
Metric fusion is another widely used approach of hybrid sampling. It directly combines different informativeness metrics. For example, one could directly sum up all metrics and select the samples with the highest values for annotation. Ranked batch-mode (Cardoso et al., 2017) can adaptively fuse multiple metrics in AL.
@@ 3.3.4 Learnable Sampling
One of the most common methods of learnable sampling is to formulate sample selection as a reinforcement learning (RL) problem, where the learner and the dataset are considered the environment, and the neural selector serves as the agent. The agent interacts with the environment by selecting a limited number of samples for annotation, and the environment returns a reward to train the neural selector. Haußmann et al. (2019) adopted a probabilistic policy network as the neural selector. The rewards returned by the environment encouraged the neural selector to choose diverse and representative samples. The neural selector is trained using the REINFORCE algorithm (Williams, 1992). In pedestrian re-identification, Liu et al. (2019) used the annotation uncertainty as the reward for training the neural selector. Agarwal et al. (2020) utilized the proposed contextual diversity as RL rewards and trains a bidirectional long short-term memory network as the neural selector. In pose estimation, Gong et al. (2022) adopted multiple agents for sample selection and directly used the performance improvement of the pose estimator as the reward for training these agents. In medical image classification, Wang et al. (2020a) employed an actor-critic framework where the critic network is used to evaluate the quality of the samples selected by the neural selector. This method has performed excellently in lung CT disease classification and diabetic retinopathy classification of fundus images.
For more works on learnable sampling in AL, such as formulating AL as few-shot learning or training neural selectors by meta-learning, please refer to the survey of Liu et al. (2022b).
@@ 4 Integration of Active Learning and Other Label-Efficient Techniques
Various methods have been proposed to reduce the large amount of labeled data required for training deep models, such as active learning, semi-supervised learning, self-supervised learning, etc. These methods are collectively called label-efficient deep learning (Jin et al., 2023a). Label-efficient learning is a broad concept that includes all related technologies designed to improve annotation efficiency. In §3, we summarized the core methods in AL, including the evaluation of informativeness and sampling strategies. However, there is still room for AL to further improve the label efficiency. For example, AL has not used unlabeled data for training, has not considered distribution shift, and still needs to annotate the whole image in fine-grained tasks such as segmentation. Integrating active learning with other label-efficient techniques can increase annotation efficiency. While some efforts have been made, existing surveys have not yet systematically organized and categorized this line of work. Hence, as one of the main contributions of this survey, we comprehensively reviewed the integration of AL with other label-efficient techniques, including semi-supervised learning, self-supervised learning, domain adaptation, region-based annotation, and generative models. Additionally, how each surveyed work integrated with other label-efficient techniques is summarized in Table 4.
@@
[…]
Zhou et al. (2017) 2017 CVPR Single Model Data Disagreement Entropy KL Divergence - - Hybrid - Fusion
@@
Yang et al. (2017) 2017 MICCAI Model Disagreement Variance Cover-based Cosine Similarity Hybrid - Multi-round
@@
Mackowiak et al. (2018) 2018 BMVC Model Disagreement Vote Entropy - - Top-k Patch
@@
Sourati et al. (2018) 2018 DLMIA Gradient-based Metrics Fisher Information - - Diversity - Solve Programming Problem
@@
Kuo et al. (2018) 2018 MICCAI Model Disagreement JS Divergence - - Diversity - Solve Programming Problem
[…]
@@
[…]
@@
Liu et al. (2019) 2019 ICCV - - - - Learnable - Reinforcement Learning
Aghdam et al. (2019) 2019 ICCV Data Disagreement BALD - - Top-K
[…]
@@
Sourati et al. (2019) 2019 TMI Gradient-based Metrics Fisher Information - - Diversity - Solve Programming Problem
@@
[…]
Siddiqui et al. (2020) 2020 CVPR Multiple Inferences - MC Dropout Data Disagreement Entropy KL Divergence - - Hybrid - Fusion Superpixel
@@
Gao et al. (2020) 2020 ECCV Data Disagreement Variance - - Top-k Consistency
Wang et al. (2020b) 2020 ECCV - - Discrepancy-based H-Divergence Learnable
@@
Lin et al. (2020) 2020 ECCV - - Clustering L2 Distance Diversity - Clustering
Ash et al. (2020) 2020 ICLR Gradient-based Metrics Gradient - - Diversity - Clustering
@@
Dai et al. (2020) 2020 MICCAI Gradient-based Metrics Gradient - - Latent Space Optimization & Nearest Neighbour Search Slice
@@
Wang et al. (2020a) 2020 MICCAI - - - - Learnable - Reinforcement Learning
@@
Huang et al. (2020) 2020 TMI Model Disagreement Hausdorff Distance - - Top-k
[…]
@@
[…]
@@
[…]
@@
[…]
Wu et al. (2021a) 2021 ICCV Single Model Entropy Clustering Color Difference Surface Variation Diversity - Clustering Hybrid - Fusion Superpixel
[…]
@@
Peng et al. (2021) 2021 ICCV Model Disagreement L2 Distance Cover-based Cardinal of Difference Set Hybrid - Fusion
Liu et al. (2021b) 2021 ICCV Gradient-based Metrics Influence - - Top-k
@@
Zhou et al. (2021b) 2021 MedIA Single Model Data Disagreement Entropy KL Divergence - - Hybrid - Fusion
Wu et al. (2021b) 2021 MedIA Performance Estimation - Learnable Data Disagreement Loss KL Divergence - - Hybrid - Fusion
Zhou et al. (2021a) 2021 MICCAI Performance Estimation - Learnable Dice - - Top-k
@@
[…]
@@
[…]
@@
[…]
Wang et al. (2022b) 2022 AAAI Gradient-based Metrics Gradient - - Top-k
Liu et al. (2022a) 2022 arXiv Multiple Inferences - Data Augmentations Variance - - Top-k Pseudo-label Superpixel
Gong et al. (2022) 2022 CVPR - - Discrepancy-based MMD Learnable - Reinforcement Learning & Meta Learning
[…]
@@
[…]
Parvaneh et al. (2022) 2022 CVPR Data Disagreement Inequality - - Diversity - Clustering
[…]
@@
[…]
@@
[…]
Hu et al. (2022) 2022 ECCV Multiple Inferences - Data Augmentations Data Disagreement Entropy KL Divergence - - Hybrid - Multi-round Pseudo-label Superpixel
[…]
Dai et al. (2022) 2022 MedIA Gradient-based Metrics Gradient - - Latent Space Optimization & Nearest Neighbour Search Slice

### v3

@@ 3.3 Sampling Strategy
With a well-developed informativeness metric, most deep AL works simply adopted top-k to select samples with the highest informativeness for annotation. However, existing informativeness metrics face several issues, such as redundancy and class imbalance. These issues are exacerbated due to the unique characteristics of medical images. Despite the high variability, medical images of the same region-of-interest (ROI) could be classified into several groups, and images within each group share a high similarity (Zheng et al., 2019). Also, class imbalance is notorious in medical image analysis since the healthy objects often outnumber the diseased ones. Instead of proposing a better informativeness metric, we can improve the sampling strategy upon the top-k selection to effectively resolve these issues above. Besides, specific sampling strategies can also be used for combining multiple informativeness metrics. Furthermore, with the recent development of deep AL, more and more studies directly employ neural networks for sample selection. In this context, we no longer evaluate informativeness but directly choose informative samples with neural networks. Regrettably, despite the importance of sampling strategies in AL, prior works or surveys have seldom discussed their specific attributes. As one of the contributions of this survey, we systematically summarize different sampling strategies in AL, including diversity sampling, class-balanced sampling, hybrid sampling, and learnable sampling. The taxonomy of different sampling strategies in AL is shown in Fig. 5.
@@ 3.3.1 Diversity Sampling
Diversity strategies aim to reduce sampling redundancy in active learning, meaning certain selected samples are highly similar to each other. The lack of diversity leads to the waste of the annotation budget. Besides, redundancy in the training set causes the deep models to overfit to limited training samples, thus leading to a performance drop. Therefore, many AL methods employ diversity sampling to mitigate the redundancy in selected samples. In this section, we discuss four strategies of diversity sampling, including clustering, farthest-first traversal, determinantal point process (DPP), and specific strategies tailored to certain informativeness metrics.
@@
Clustering is one of the most commonly used strategies of diversity sampling. This strategy improves the coverage of the entire feature space, thereby easily boosting diversity. Ash et al. (2020) employed k-Means++ clustering on gradient embeddings to select diverse uncertain samples. Besides, Citovsky et al. (2021) boosted margin-based uncertainty sampling with hierarchical clustering. They selected samples with the smallest margins within each cluster. When the number of queries exceeded the number of clusters, samples from smaller clusters were prioritized. This method can scale to a huge annotation budget (e.g., one million). Zheng et al. (2019) incorporated hierarchical clustering with cover-based AL. In their experiments, clustering showed consistent performance improvement in multiple medical imaging datasets, which demonstrated that clustering does improve the sampling diversity.
It is important to highlight that clustering in this section is different from that of in §3.2.1. To ensure that the selected samples are sufficiently representative, clustering-based AL generally chooses samples that are closest to the cluster centers. However, when clustering is used to enhance diversity, we can not only select samples closest to the cluster centers, but also select samples with the highest uncertainty, or even randomly select within each cluster. Therefore, clustering can serve as a plug-and-play technique to conveniently enhance the sampling diversity in AL.
@@
Determinantal point process is a stochastic probability model for selecting subsets from a larger set. DPP reduces the probability of sampling similar elements to ensure diversity in the results. Bıyık et al. (2019) employed two DPPs for sample selection: Uncertainty DPP is based on uncertainty scores, while Exploration DPP aims to find samples near decision boundaries. Then, sampling results from both DPPs were sent for expert annotation. However, DPP is more computationally intensive compared to clustering. Ash et al. (2020) compared the performance and time cost of using k-Means++ and k-DPP. Results showed that their performance is similar, but the time cost for k-Means++ is significantly lower than that for k-DPP. Besides, Mi et al. (2020) adopted DPP in AL for medical image reconstruction, please refer to §5.3 for details.
Random Selection could also be used for better diversity. In prostate segmentation of MRI, Gaillochet et al. (2023a) randomly partitioned the entire dataset into different batches, which were referred to as ‘stochastic batches’. Batches with the highest uncertainty scores were selected for annotation. Experimental results showed that the stochastic batches consistently improved the performance of various uncertainty-based AL methods under an extremely low budget. Their extended works (Gaillochet et al., 2023b) further illustrated the effectiveness of stochastic batches on the anterior and posterior hippocampus segmentation.
Farthest-first Traversal is also a widely used strategy for diverse queries, which was first adopted by Sener and Savarese (2018). This strategy requires the distance between sampling points to be as large as possible, which leads to a more uniform distribution of selected samples in the feature space. Li et al. (2023) adopted a farthest-first traversal strategy with cosine distance for a diverse initial labeled dataset. Experiments on breast ultrasound, liver CT, and chest X-ray segmentation showed significant effectiveness of the farthest-first traversal. Besides, Agarwal et al. (2020) and Caramalau et al. (2021) improved the diversity with farthest-first traversal with their proposed contextual diversity and GNN-augmented features, respectively.
Other strategies: In uncertainty-based AL, BatchBALD (Kirsch et al., 2019) extended BALD-based uncertainty AL to batch mode. Results showed that BatchBALD improved the sampling diversity compared to (Gal et al., 2017). FI-based methods formulated AL as a semi-definite programming (SDP) problem to improve sampling diversity and various methods were employed for solving SDP. Sourati et al. (2019) used a commercial solver to solve SDP, while Ash et al. (2021) proposed a greedy algorithm to adapt to high-dimensional feature space. In skin lesion analysis, Shi et al. (2019) introduced image hashing for diversity sampling. In their proposed method, the first principal component of each image was used for feature representation. Then they mapped similar images into the same buckets using local sensitivity hashing. Samples were uniformly selected from each bucket for human annotation.
@@ 3.3.2 Class-balance Sampling
Class imbalance is a common issue for DL in medical image analysis, where a small set of classes have many samples while the others only contain a few samples (Zhang et al., 2023b). For example, long tail distribution of classes existed in almost all tasks of medical image classification, such as skin lesion classification and whole-slide image classification. Training on imbalanced datasets can lead to the overfitting of the majority classes and underfitting of the minority classes. Apart from dealing with class imbalance during training, AL mitigates class imbalance by avoiding over-annotation of the majority classes and enhancing the annotation of the minority classes during dataset construction.
@@
Classification: In a class-imbalanced COVID-19 dataset, Chong et al. (2021) evaluated multiple informativeness scores and sampling strategies. Results showed that diversity sampling is more favored for class-imbalance. Jin et al. (2022c) assumed that samples closer to the tail of the distribution are more likely to belong to the minority classes. Thus, the tail probability is equivalent to the likelihood of minority classes. Specifically, they trained a VAE for feature extraction and adopted copula to estimate the tail probabilities upon VAE features. Finally, informative samples were selected with clustering and unequal probability sampling. The proposed method was validated on the ISIC 2020 dataset, which has a long-tailed distribution. Kothawade et al. (2022c) used submodular mutual information to focus more on samples of minority classes. They achieved excellent results on medical classification datasets in five different modalities, including X-rays, pathology, and dermoscopy. In blood cell detection under microscopy, Sadafi et al. (2019) requested expert annotation of a sample whenever its classification probability of the minority class exceeded 0.2. Besides, Choi et al. (2021b) directly estimated the probability of a classifier making a mistake for a given sample and decomposed it into three terms using Bayesian rules. First, they trained a VAE to estimate the likelihood of the data given a predicted class. Then, an additional classifier was trained upon VAE features to estimate class prior probabilities and the probability of mislabeling a specific class. By considering all three probabilities, they successfully mitigated class imbalance in AL. The proposed method achieved good performance on stepwise class-imbalanced CIFAR-10 and CIFAR-100 datasets. For uncertainty-based methods, Bengar et al. (2022) introduced an optimization framework to maintain class balance. They compensated the query of minority classes with the most confident samples of that class, leading to a more balanced class distribution in the queried dataset.
@@
Segmentation: Due to certain AL methods selecting regions instead of the entire image for annotation, there is a need to ensure that the selected regions contain rare or small objects (e.g., optic chiasma or optic nerve in head and neck multi-organ segmentation). Cai et al. (2021) and Wu et al. (2022b) both proposed class-balanced sampling strategies for such scenarios, as detailed in §4.3.
@@ 3.3.3 Hybrid Sampling
In AL, more and more works use multiple informativeness metrics simultaneously. However, how to effectively integrate multiple metrics remains a critical issue. This issue is addressed by the hybrid sampling discussed in this section. Two approaches to hybrid sampling are often used, including multi-round sampling and metric fusion.
Multi-round sampling first selects a subset of samples based on one particular informativeness metric and continues sample selection within this subset based on another informativeness metric. Multi-round sampling is widely used in AL for medical image analysis for its convenience (Shen et al., 2020; Li and Yin, 2020; Wang and Yin, 2021). For example, SA (Yang et al., 2017) performed representativeness sampling based on uncertainty to reduce redundancy in the sampled set. Besides, Wu et al. (2022b) employed an adaptive strategy that sets dynamic weights to adjust the budget of representativeness and uncertainty sampling. The weight of representativeness sampling is larger initially, while the situation is reversed in the latter phase. This is because representativeness methods can quickly spot typical data, while uncertainty methods continuously improve the model by querying samples with erroneous predictions.
Metrics fusion is another widely used approach of hybrid sampling. It directly combines different informativeness metrics. For example, one could directly sum up all metrics and select the samples with the highest values for annotation. Metrics fusion is also widely used in AL of the medical domain (Li et al., 2024; Zhou et al., 2021c; Wu et al., 2021). Besides, ranked batch-mode (Cardoso et al., 2017) can adaptively fuse multiple metrics in AL.
@@ 3.3.4 Learnable Sampling
One of the most common methods of learnable sampling is to formulate sample selection as a reinforcement learning (RL) problem, where the learner and the dataset are considered the environment, and the neural selector serves as the agent. The agent interacts with the environment by selecting a limited number of samples for annotation, and the environment returns a reward to train the neural selector. In medical image classification, Wang et al. (2020b) employed an actor-critic framework where the critic network is used to evaluate the quality of the samples selected by the neural selector. This method has performed excellently in lung CT disease classification and diabetic retinopathy classification of fundus images. Besides, Haußmann et al. (2019) adopted a probabilistic policy network as the neural selector. The rewards returned by the environment encouraged the neural selector to choose diverse and representative samples. The neural selector is trained using the REINFORCE algorithm (Williams, 1992). Agarwal et al. (2020) utilized contextual diversity as RL rewards and trains a bidirectional long short-term memory network as the neural selector.
For more works on learnable sampling in AL, such as formulating AL as few-shot learning or training neural selectors by meta-learning, please refer to the survey of Liu et al. (2022).
@@ 4 Integration of Active Learning and Other Label-Efficient Techniques
As discussed in §1, the high annotation cost has severely dragged down the development of DL in medical image analysis. Despite the wide use of AL in medical image analysis, various methods have been proposed to reduce the large amount of labeled data required for training deep models, such as semi-supervised and self-supervised learning, etc. These methods, including active learning, are collectively called label-efficient deep learning (Jin et al., 2023a). Label-efficient learning is a broad concept that includes all related technologies designed to improve annotation efficiency. In the real-world practice of AL in medical image analysis, there is still room for higher label efficiency by integrating AL with other label-efficient techniques. For the example of AL in medical image segmentation, since many samples were left unlabeled in the cycle of AL, we could further include them to achieve better performance by integrating AL with semi-supervised learning. The rapid development of self-supervised learning in medical image analysis introduced many powerful pre-trained models (Taleb et al., 2020). These models are also valuable in AL of medical image analysis for their superior ability of feature extraction. For another circumstance, since the ROI in medical imaging is usually small, we could select and annotate the informative regions that contain the ROI in AL instead of annotating the whole image. As a result, integrating active learning with other label-efficient techniques holds significant potential to increase annotation efficiency. However, existing surveys have not yet systematically organized and categorized this line of research. Hence, as one of the main contributions of this survey, we comprehensively reviewed the integration of AL with other label-efficient techniques, including semi-supervised learning, self-supervised learning, domain adaptation, region-based annotation, and generative models. Additionally, how each surveyed work integrated with other label-efficient techniques is summarized in Table 4.
@@
Zhu and Bento (2017) 2017 arXiv Single Model Distance to Decision Boundary - - Top-k
Zhou et al. (2017) 2017 CVPR Single Model Multiple Inferences - Data Disagreement Entropy KL Divergence - - Hybrid - Fusion
@@
Yang et al. (2017) 2017 MICCAI Multiple Inferences - Model Disagreement Variance Cover-based Cosine Similarity Hybrid - Multi-round
@@
Mackowiak et al. (2018) 2018 BMVC Multiple Inferences - Model Disagreement Vote Entropy - - Top-k Patch
@@
Sourati et al. (2018) 2018 DLMIA Gradient-based Uncertainty Fisher Information - - Diversity - Solve Programming Problem
@@
Kuo et al. (2018) 2018 MICCAI Multiple Inferences - Model Disagreement JS Divergence - - Diversity - Solve Programming Problem
Mahapatra et al. (2018) 2018 MICCAI Multiple Inferences - MC Dropout Variance - - Top-k
@@
Zheng et al. (2019) 2019 AAAI - - Cover-based Cosine Similarity Diversity - Clustering
@@
Tran et al. (2019) 2019 ICML Multiple Inferences - MC Dropout BALD - - Top-k
@@
Sourati et al. (2019) 2019 TMI Gradient-based Uncertainty Fisher Information - - Diversity - Solve Programming Problem
@@
Shui et al. (2020) 2020 AISTATS Single Model Entropy, Least Confidence Discrepancy-based Wasserstein Distance Hybrid - Fusion
Siddiqui et al. (2020) 2020 CVPR Multiple Inferences - MC Dropout Multiple Inferences - Data Disagreement Entropy KL Divergence - - Hybrid - Fusion Superpixel
@@
Gao et al. (2020) 2020 ECCV Multiple Inferences - Data Disagreement Variance - - Top-k Consistency
Wang et al. (2020c) 2020 ECCV - - Discrepancy-based H-Divergence Learnable
@@
Lin et al. (2020) 2020 ECCV - - Clustering-based L2 Distance Diversity - Clustering
Ash et al. (2020) 2020 ICLR Gradient-based Uncertainty Gradient - - Diversity - Clustering
@@
Dai et al. (2020) 2020 MICCAI Gradient-based Uncertainty Gradient - - Latent Space Optimization & Nearest Neighbour Search Slice
@@
Wang et al. (2020b) 2020 MICCAI - - - - Learnable - Reinforcement Learning
@@
Huang et al. (2020) 2020 TMI Multiple Inferences - Model Disagreement Hausdorff Distance - - Top-k
Su et al. (2020) 2020 WACV Single Model Entropy Discrepancy-based H-Divergence Hybrid - Fusion
@@
Fu et al. (2021) 2021 CVPR Adversarial Training Disagreement of Classifiers, margin Discrepancy-based H-Divergence Hybrid - Fusion
@@
Prabhu et al. (2021) 2021 ICCV Single Model Entropy - - Diversity - Clustering
@@
Du et al. (2021) 2021 ICCV - - Discrepancy-based Semantic and distinctive scores Hybrid - Fusion
Shin et al. (2021) 2021 ICCV Multiple Inferences - Model Disagreement Inequality - - Diversity - Clustering Pseudo-label Pixel, Point
Rangwani et al. (2021) 2021 ICCV Adversarial Samples KL Divergence Cover-based - Submodular KL Divergence Bhattacharya Coefficient Hybrid - Fusion
@@
Liu et al. (2021) 2021 ICCV Gradient-based Uncertainty Influence - - Top-k
@@
Zhou et al. (2021c) 2021 MedIA Single Model Multiple Inferences - Data Disagreement Entropy KL Divergence - - Hybrid - Fusion
Wu et al. (2021) 2021 MedIA Performance Estimation - Learnable Multiple Inferences - Data Disagreement Loss KL Divergence - - Hybrid - Fusion
Zhou et al. (2021b) 2021 MICCAI Performance Estimation - Learnable Dice - - Top-k
@@
Ash et al. (2021) 2021 NeurIPS Gradient-based Uncertainty Fisher Information - - Diversity - Solve Programming Problem
@@
Chen et al. (2021) 2021 TPAMI Single Model (in Feature Space) Entropy - - Top-k
@@
Xie et al. (2022b) 2022 AAAI Single Model Margin Density-based Energy Hybrid - Multi-round
Wang et al. (2022b) 2022 AAAI Gradient-based Uncertainty Gradient - - Top-k
Xie et al. (2022d) 2022 CVPR Single Model Margin, Gradient - - Top-k
@@
Zhang et al. (2022b) 2022 CVPR Single Model Entropy - - Top-k
Parvaneh et al. (2022) 2022 CVPR Multiple Inferences - Data Disagreement Inequality - - Diversity - Clustering
Xie et al. (2022a) 2022 CVPR Single Model Entropy - - Top-k Patch
@@
Wang et al. (2022c) 2022 ECCV - - Density-based KNN Density Diversity - Clustering w/ Regularization Consistency
@@
Chen et al. (2022b) 2022 ECCV Gradient-based Uncertainty Gradient - - Top-k
Hwang et al. (2022) 2022 ECCV Single Model Margin Discrepancy-based MMD Hybrid - Multi-round Pseudo-label
Yi et al. (2022) 2022 ECCV Single Model Least Confidence Self-supervised Learning Loss of Pretext Task Hybrid - Multi-round
Wu et al. (2022b) 2022 ECCV Single Model Entropy Density-based GMM Hybrid - Multi-round Superpixel
Mahmood et al. (2022) 2022 ICLR - - Discrepancy-based Wasserstein Distance Diversity - Solve Programming Problem
Hacohen et al. (2022) 2022 ICML - - Density-based Inverse Average Distance to KNN samples Diversity - Clustering
Jin et al. (2022a) 2022 Information Sciences - - Clustering-based Cosine Similarity Diversity - Clustering
Jin et al. (2022c) 2022 KBS - - Clustering-based L2 Distance Class-balance
Jin et al. (2022b) 2022 KBS - - Clustering-based L2 Distance Diversity - Farthest-first Traversal
Dai et al. (2022) 2022 MedIA Gradient-based Uncertainty Gradient - - Latent Space Optimization & Nearest Neighbour Search Slice